10 years ago, a 20 year old boy told me he is 10 years older than me. how old i am now?
Answer by mikeebsc(26)

Beginning word problems can be very hard to understand, but once you get it, you've got it for good, like riding a bike.
In this problem, we see the words "10 years ago", so we know we must subtract 10 from something. Since we do not know what that something is, we represent it with the variable x, so we get x-10. Now, we also know at this same point in time (the reason for it being equal to) a 20 yr old boy said he was 10 years older which translates into the language of math as 20+10. You can look at it like this, all of this happened at the same time, therefore they are equal to each other; you are simply translating the language English to the language of mathematics. Trying to explain this is really the hardest part, as i do not know what level or grade you are in, so I am trying multiple ways. We can also say that there are 2 parts to this sentence separated by a comma, but the events happened at the same time, therefore they are equal; hence the equal sign.
Form the equation to solve:
x-10=20+10
add 10 to both sides to get x alone (isolate the variable)
+10 = +10
x=20+10+10 (The variable x is isolated, so we may simply add the numbers together on the other side)
x=40
You are 40 now :)
Hope this helps!!
If I can help any further, please let me know!